OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of 2x2 + 4 = 0

OpenStudy (anonymous):

discriminant= b^2-4ac a=2 b=0 c=4 discriminant= 0- 4(4)(2) The discriminant is -32 which is less than zero

OpenStudy (anonymous):

so no real root

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of 8x2 = 0 thats no real root two?